Find three consecutive even numbers whose sum is 1368.
small number =
middle number =
larger number =

Answer by ikleyn(52778)

.

Let x be the number in the middle.

Then two other numbers are (x-2) and (x+2).


The sum of the three numbers is 

    (x-2) + x + (x+2) = 1368,

or

    3x = 1368

     x = 1368/3 = 456.


ANSWER.  The numbers are 454, 456 and 458.

Solved.

----------------

In the future,  if you have such problem with different sum,  simply divide
the sum by  3:  it will give you the middle number.

This simple trick works  ALWAYS,  saving your time.


//////////////////


To see many other similar solved problems,  look into the lesson
    - Problems with consecutive integer numbers; odd/even consecutive integer numbers
in this site.

Learn the entire subject from there.